OpenStudy (anonymous):

Anyone can help me with this problemIf Vector A^ = (7.00i -9.00J ) units, VectorB^ = (-9.00i + 4.00J ) units, and Vector C^ = (25.0i + 20.0J ) units, determine a and b such that aA^ + bB^ +C^ = 0. a =? b =?

OpenStudy (anonymous):

for A^ = (7.00i -9.00J ) |A| = sqrt(49 + 81) = sqrt(130) for B^ = (-9.00i + 4.00J ) |B| sqrt (81 + 16) = sqrt(97) for C^ = (25.0i + 20.0J ) |C| = sqrt(625+400) = sqrt (1025) a sqrt(130) + b sqrt(97) + sqrt (1025) = 0 then you will get a value of a and b

OpenStudy (anonymous):

I did it as you showed it on above, but i got the answer wrong. How would I find a and b?

OpenStudy (anonymous):

you got it???

OpenStudy (anonymous):

I dont know how to find for a and b?

OpenStudy (anonymous):

Would you tell me how do I find?

OpenStudy (anonymous):

Now, I got it!

OpenStudy (anonymous):

Thanks for your help?

OpenStudy (anonymous):

congrats! welcome . two thumbs for you! ^^

OpenStudy (anonymous):

Did u get the answer for a and b too? What did you get it if you got the answer?

OpenStudy (anonymous):

yes...., do you know?